1.how do i write the exponets of
x*x*y*y*y?
is it x2 y3?
2.how do i write the exponets of
a/b*a/b*c/d*c/d*c/d?

Answer by checkley71(8403)

x*x*y*y*y=x^2y^3 Correct!
a/b*a/b*c/d*c/d*c/d=(a/b)^2(c/d)^3 This is exactly the same problem as the x*y problem only with fractions. Don't let fractions confuse you. Treat them like any other term or number.